we are having trouble with this equation:
8(x+3) = (8x-12)

Answer by checkley77(12844)

8(x+3) = (8x-12)
8X+24=8X-12
8X-8X=-12-24
0X=-36 NO UNIQUE SOLUTION FOR X.